Calculate the difference in electronegativity for a bond between C and F?
Calculate the difference in electronegativity for a bond between B and Te?
Calculate the difference in electronegativity for a bond between N and CI?
Which bonds (A,B,C) would be considered to be polar
1
Expert's answer
2020-11-17T10:32:23-0500

A) C-F= 3.98 - 2.55 = 1.43.

B) B-Te= 2.1 - 2.04 = 0.6.

C) N-Cl= 3.16 - 3.04 = 0.12.

In chemistrypolarity is a separation of electric charge leading to a molecule or its chemical groups having an electric dipole moment, with a negatively charged end and a positively charged end.

A.